The closing price of Schnur Sporting Goods Inc. common stock is uniformly distributed between $18 and $31 per share.

What is the probability that the stock price will be:
a. More than $24? (Round your answer to 4 decimal places.)
  Probability    
b. Less than or equal to $20? (Round your answer to 4 decimal places.)
  Probability    
Answer #1

X ~ uniform( a,b) = Uniform ( 18 , 31)

Where a = 18 , b = 31

a)

P( X > x) = (b - x) / (b - a)

So,

P( X >24 ) = (31 - 24) / ( 31 - 18)

= 0.5385

b)

P( X <= x) = (x - a) / (b - a)

P( X <= 20) = (20 - 18) / (31 - 18)

= 0.1538